OpenStudy (anonymous):

why i^2 = -1, where ' i ' is (iota).

OpenStudy (anonymous):

\[i = \sqrt{-1}\]

OpenStudy (anonymous):

yes then??

OpenStudy (anonymous):

\[i=\sqrt{-1}\] \[i^2 = (\sqrt{-1})^2 = -1\]

OpenStudy (shubhamsrg):

its the definition of iota.. that i = sqrt(-1) .. it'd be like asking why 2 is called two ? its the definition,,no theoritical reason..

OpenStudy (anonymous):

yes but what about this one

OpenStudy (anonymous):